a bus starting from one point to another point in 45 minutes with the speed of 40 km/hr. if the speed increased by 10 km/hr how much time it will take to cover the same distance. wiki

Answer by Alan3354(69443)

a bus starting from one point to another point in 45 minutes with the speed of 40 km/hr. if the speed increased by 10 km/hr how much time it will take to cover the same distance. wiki
-----------
45 mins = 0.75 hours
Find the distance
d = r*t
------
Find the time at 50 km/hr
d = r*t
t = d/r